Let X be random variable with the binomial distribution with parameters n and 0 < p < 1.

(1) Show that (P(X = x) / P(X = x -1)) - 1 = np + (p - x)) / (x(1-p)) for any 1 ≤ x ≤ n.

(2) Show that when 0 ≤ x < (n + 1)p , P(X = x) is an increasing function x and for (n + 1)p < x ≤ n, P(X = x) is a decreasing function x.
Answer #1

(1)

(P(X = /P(X = 1 - 1)) - 1

- (3) *1 – p)=16*. 12–14 - p)=+

=(n!/x!(n − 2)!p/(n!/(1 - 1)!(n - 1 + 1)!)(1 - p) - 1

= (n - 1+1)p/c(1 - p) – 1

(mp – Tp+p)/(– Tp) - 1 = ((mp – Tp+p) - ( - Tp))/(x - Tp)

(app – x))/3(1 -p)

(2)

For 0 ≤ x < (n + 1)p

(n + 1)p - x > 0

then,

(P(X = x) / P(X = x -1)) - 1 = np + (p - x)) / (x(1-p)) = ((n + 1)p - x ) / (x(1-p)) > 0

=> (P(X = x) / P(X = x -1)) > 1

=> P(X = x) > P(X = x -1)

=> P(X = x) is an increasing function of x

For (n + 1)p < x ≤ n

=> (n + 1)p - x < 0

(P(X = x) / P(X = x -1)) - 1 = np + (p - x)) / (x(1-p)) = ((n + 1)p - x ) / (x(1-p)) < 0

=> (P(X = x) / P(X = x -1)) < 1

=> P(X = x) < P(X = x -1)

=> P(X = x) is an decreasing function of x